Perform cakculations based on the following financial statements Round all answers to the nearest tenth unless otherwise specified a) What is the debt ratio in 2018? In 2018, the company has a lower proportion of debt to assets than it did in 2017 O False b) What is tha quick (acid-test) ratio in 2018? Which of the following is true? Mashayekhi and Meckstroth Manufacturing, Ind. Consolidated Balance Sheets as of June 30, n 2017? fin thousonds ASSETS Current assets: 2018 2017 Cash and equivalents Short-term investments Accounts receivable Inventories Prepaid insurance 5 2,140 $ 4,000 980 7,278 4,195 910 90 11,757 10,175 22,330 22,330 (8801 21,340 21,450 1,485 770 A. A lower quick ratio indicates a greater ability to pay off current liabilities. OB. A higher quick ratio indicates a greater ability to pay off curent labilities. c) What is the inventory tunover ratio in 2013? d) What is the cost of goods sold in 20187 (Round to the nearest dollar.) S e) What is the gross margin in 2018? ) What is the tax rate? g) What is the return on total assets in 2018? h) What is the return on common equity in 2018? Total current assets Property, plantand equipment Accumulated deprecietion -PP&E 990 Net PP&E LIABILITIES AND OWNERS EQUITY Current liabilities Accounts payable Notes payable Wages payable Income tax payable 1,325 970 460 550 Total current liabilities Long-term debt Other llabilities 1,900 1,455 9,860 13,460 6,800 $20,360 $21,715 8,600 OTAL LIABILITIES Owners equity: Comman stock- 2000 shares outstanding Capital in excess of par value Retained earnings OTAL OWNERS EQUITY 20 20 80 537 9.810 $12,737 9,910 TOTAL LIABILITES AND OWNREQUTY$33,097 $31,625 Mashayekhi and Meckstroth Manufacturing, Ind. Consolidated Income Statements as of the year ending lune 30, fin thousands) Sales revenue 2018 2017 $20,110 $16,410 7,500 7,100 408 410 11,352 8,492 6,200 7,800 Direct materials Direct labor 408 850 Gross profit Earnings before interest and taxes Income before income taxes Selling, general and administrative expense 5,152 Interest expense 4,712 1,885 Income tax expense 15) NET INCOME (LOSS)
Answer #1

(a)

(i) Debt ratio = Long-term debt / Total assets

2018: 9,860 / 33,097 = 0.3

2017: 13,460 / 31,625 = 0.4

(ii) True

In 2018, debt ratio was lower than in 2017 (0.3 < 0.4).

(b)

(i) Quick ratio = (Cash & equivalent + Short term investment + Accounts receivable) / Current liabilities

2018: (2,140 + 1,485 + 7,278) / 1,900 = 10,903 / 1,900 = 5.7

(ii) Option (B)

The higher (lower) the quick ratio, the higher (lower) the ability to pay off current liabilities.

(c) Inventory turnover ratio = Cost of goods sold / Average Inventory = (Sales revenue - Gross profit) / Average inventory

= (20,110 - 11,352) / [(770 + 910) / 2] = (8,758 x 2) / 1,680 = 10.4

(d) Cost of goods sold = Sales revenue - Gross profit = 20,110 - 11,352 = 8,758

NOTE: As per Answering Policy, 1st 4 parts are answered.